# Generalizing the SINDy approach with nested neural networks

### v1

@@ 3. The Nested SINDy approach
where $\mathcal{N}$ denotes our nested model, parameterized by $\theta$. Consequently, the resolution of this nonlinear optimization problem may be significantly more challenging and algorithms that work well in the linear case are not necessarily convergent in the nonlinear case.
The method then primarily depends on the choice of the architecture $\mathcal{N}$. The goal is to introduce new layers that achieve favorable trade-offs between expressivity and optimization complexity. In this work, we propose two architectures, which are described in the following sections: the PR model (in Section 3.1) and the PRP model (in Section 3.2). From now on, we call the basic SINDy layer, given by a projection onto $L(\mathcal{F})$ (2.2), the radial layer.
3.1. The PR Model
The PR (Polynomial-Radial) model augments the basic SINDy framework by introducing a polynomial layer that operates before the usual SINDy radial layer. This layer constructs a variety of monomials from the input variable $x$, represented as follows:
@@ 3.1. The PR Model
with $\omega_{0,0}$ acting as the constant term of the polynomial, effectively substituting the bias. Another choice is to limit the sum over $i+j$ to a maximum value $d$ to reduce the number of terms in the polynomial, and thus obtain bivariate polynomials up to degree $d$.
As an example, for one input variable, if =$\mathcal{F}=\{\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits,\cos\}$ and $d=2$, then the PR model can learn all functions with expression
@@
This PR layer can also be seen as a pure polynomial layer combined with a linear layer. We, hence, consider the PR model to have four layers: a polynomial layer, a linear layer, a radial layer, and a final linear layer. The last two layers are identical to the standard SINDy model, while the first two layers are new additions. Figure 1 illustrates this structure.
@@
$f_{\theta,\text{PR}}(x)=\sum\limits_{j=1}^{l}c_{j}f_{j}\left(\sum\limits_{i=0}^{d}\omega_{i,j}x^{i}\right)+B,$ (3.1)
where $\theta$ includes all the trainable parameters of the model. The functions $f_{j}$ are derived from the specified function set $\mathcal{F}$, while $\omega_{i,j}$ are the weights of the polynomial layer, and $c_{j}$ and $B$ are the weights and bias of the final linear layer, respectively. The parameters $\omega_{i,j}$ now depend on $j$ since they are in the $j$-th radial layer. It is important to note that the radial and polynomial layers are not associated with any adjustable parameters.
@@ 3.2. The PRP Model
The PRP (Polynomial-Radial-Polynomial) model enhances the PR model by introducing an additional polynomial layer following the radial layer. This structure significantly improves the ability of the model to represent complex interactions within datasets. For example, the PRP model can express the function $f(x)=\arctan(x)\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x)$, (assuming arctangent and sine are in the dictionary), while the PR model cannot, as it is not a linear combination of the functions contained within the dictionary. In the PRP model, the outputs of the radial layer are first processed through a linear layer. This linear layer, typically fixed in size (set to 2 in our experiments), serves as an intermediate stage, transforming the outputs of the radial layer into a new set of variables. These variables are then fed into a subsequent polynomial layer, which allows for the formation of various monomial combinations of the outputs of the linear layers, such as squaring or multiplying them.
@@
$f_{\theta,\text{PRP}}(x)=\sum\limits_{1\leq\lvert i\rvert\leq d}\omega_{i_{1},i_{2},\dots,i_{l}}^{\text{PR}}f_{\theta_{1},\text{PR}}(x)^{i_{1}}f_{\theta_{2},\text{PR}}(x)^{i_{2}}\dots f_{\theta_{l},\text{PR}}(x)^{i_{l}}+B^{\prime},$ (3.2)
where $\lvert i\rvert=i_{1}+i_{2}+\ldots+i_{l}$ is the length of the multi-index $(i_{1},\ldots,i_{l})$, $\theta$ includes all the trainable parameters of the model, $\omega_{i_{1},i_{2},\dots,i_{l}}^{\text{PR}}$ are the weights of the final linear layer, $B^{\prime}$ is the bias of the final linear layer, $l$ is the size of the output chosen for the intermediate linear layer (set to 2 in our experiments), and $f_{\theta_{1},\text{PR}}(x)$, $f_{\theta_{2},\text{PR}}(x)$, …, $f_{\theta_{l},\text{PR}}(x)$ correspond to the output of the PR model given in (3.1). The coefficients $\theta_{1},\theta_{2},\dots,\theta_{l}$ correspond to the parameters of the PR models, which are the same as $\theta$, except for the weights of the final linear layer ($c_{j}$ and $B$ in (3.1)). Figure 2 shows a graphical representation of this model.
@@
Figure 2. Structure of the PRP model for $l=2$, $d=2$, and 2 input dimensions.
@@ 4.1. Adding a regularization term to enforce sparsity
$\lambda(\text{epoch})=\lambda_{0}(1+0.4\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(\text{epoch}/10)),$
where $\lambda_{0}$ is the initial Lasso coefficient, and where the sine function is used to make the coefficient oscillate between $0.6\lambda_{0}$ and $1.4\lambda_{0}$.
@@ 4.4. Adding noise to the gradient of the loss function
The noise amplitude has to be well-tuned. It depends on the learning rate $lr$ of the optimizer, the values of the parameters $\theta$, and on the actual mean squared error $L_{\text{MSE}}$:
$\nabla_{\theta}L\leftarrow\nabla_{\theta}L+\varepsilon(lr,L_{\text{MSE}},\theta).$
@@ 5.1.1. Case 1: trigonometric function involving composition using the PR block
The Adam optimization algorithm is used, with a learning rate of $10^{-3}$. Lasso regularization is applied at a rate of $10^{-3}$. A batch size of 1 00010001\,0001 000, dictating 10 iterations per epoch, is used. Over the course of 1 00010001\,0001 000 epochs, the model weights undergo pruning every $30$ epochs with a threshold set to $0.05$. All computations are performed on a CPU. The function uses a patience counter to keep track of how many consecutive epochs have passed without significant improvement in the loss function. This is calculated by comparing the relative difference in loss between the current and previous epochs to a predefined threshold (here, $10^{-2}$). If the relative difference is lower than this threshold for $50$ consecutive epochs ($\texttt{patience}=50$), the training stops. In this case, the training stopped after $300$ epochs.
The learning process of the PR model with the specified parameters successfully converged, yielding a model that approximates the target function with the expression $x\mapsto-\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}-1.57)$ that captures the underlying dynamics of the target function within the considered domain, see Figure 3. The sparsity of the solution is significant, with only 3 nonzero coefficients out of the original 29 parameters, underscoring the effectiveness of the model in identifying the essential structure of the function within the given domain.
@@
(d) Epoch 180: the model has converged to a prediction matching the target function.
Figure 3. Case 1: evolution of the learned function $x\mapsto\cos(x^{2})$ over successive training epochs. Each subfigure represents the PR model’s predictions (in blue) against the target function (in orange) at epochs 10, 70, 120, and 180, showcasing the model’s progressive learning and convergence towards the target function. We observe that the first epochs learn on the interval $[0,3]$, while subsequent epochs are able to generalize.
@@ 5.1.2. Case 2: trigonometric function involving composition using the PRP block
To further illustrate the capabilities of the introduced PRP-Nested SINDy method (3.2) described in Section 3.2, we examine the same function as in case 1, namely $f:x\mapsto\cos(x^{2})$, considered over the interval $[0,3]$. This time, employing a dataset comprising 1 00010001\,0001 000 data points, the PRP-Nested SINDy model is tasked with replicating this function.
The model’s architecture begins with a polynomial (P) layer that computes monomials up to the second degree. A subsequent linear layer then transforms these two polynomial features into nine outputs, incorporating a bias term in the process. The radial (R) layer follows, applying a suite of predefined functions to each node, belonging to the following dictionary:
@@
The architecture then introduces a second polynomial (P) layer, and a final linear layer maps these five processed features to a single output, thus completing the structure of the model.
Training parameters are configured as follows: the Adam optimization algorithm is employed with a learning rate of $10^{-4}$. Lasso regularization is set at a rate of $10^{-3}$. The batch size is 1 00010001\,0001 000. Across 1 00010001\,0001 000 epochs, model weights are pruned every $30$ epochs using a threshold of $0.05$. We note that the loss function stagnates after around $300$ epochs. All computations are carried out on a CPU.
The PRP model’s learning process successfully converged to the expression $x\mapsto\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}+1.57)$. This expression effectively captures the underlying dynamics within the specified domain (Figure 4). The resulting model exhibits notable sparsity, with only 3 nonzero coefficients from the initial 31 parameters.
Figure 4. Case 2: comparison of the results from the PRP-Nested SINDy, which yielded the function $x\mapsto\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}+1.57)$, with the target function $x\mapsto\cos(x^{2})$, over the interval [0, 3].
@@ 5.1.3. Case 3: Trigonometric Function Multiplication using the PRP Block
The training parameters for learning the function $(x,y)\mapsto 2\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x)\cos(y)$ are specifically configured as follows: the Lasso regularization rate is set to $0.1$. In the uniform weight initialization for the linear layers, the standard deviation is set to $1.5$, and the mean is maintained at 00. The Adam optimization algorithm is employed with a base learning rate of $10^{-4}$, and the model is trained for 1 00010001\,0001 000 epochs. Pruning of model weights occurs every $50$ epochs with an initial pruning at $10$ epochs, using a pruning threshold of $0.01$. All computations are conducted on a CPU.
@@
with sparsity accounted for by 14 nonzero coefficients out of the original 32 parameters. A comparison of this function with the target function is illustrated in Figure 5. The MSE between the target function and its approximation is $5.69\times 10^{-2}$. Interestingly, the three terms with small coefficients in (5.1) (e.g. $0.039xy$) should not be manually removed from the formula after learning. Indeed, removing them leads to an almost fourfold increase in MSE, which gets bumped to $2.11\times 10^{-1}$.
@@ 5.1.4. Case 4: Perimeter of an ellipse
Since any rescaled ellipse remains an ellipse, we assume without loss of generality that $b=1$ for the rest of this experiment. To assess the relevance of our Nested SINDy technique, we will compare its performance against the two reference solutions: Ramanujan’s approximation, and a linear interpolation between the two endpoints of the interval $[1,30]$. Indeed, using a linear interpolation makes sense for large values of $a$, since $P$ is equivalent to $4a$ when $a\to\infty$, as can be seen by inspecting (5.1.4).
To find formulas that approximate the perimeter of an ellipse, we train two PRP models. The first one is trained on the interval $[1,5]$ and the second one on the interval $[1,25]$. We run 50 learning sessions for each model and keep the model with the shortest associated formula. In its best run, the first model converged to a basic quadratic polynomial:
@@
The second model converged to a more complex expression which is too long to be displayed here. In this second model, 25 out of the 64 parameters are nonzero, corresponding to approximately 40% sparsity. The MSE obtained at the end of the training is $1.97$. This metric, as the previous ones, is biased because the training is stopped at an arbitrary step and there is no final tuning step. Hence, we perform a final tuning step, by writing the final expression of the model and performing a gradient descent to only minimize the MSE. The later mentioned results include this final tuning step and show that the effective MSE can be significantly reduced.
@@
Overall, this experiment provides an insight into the ability of the Nested SINDy approach to discover an approximation on a classical problem without knowledge other than the data points. It demonstrates that this method can be used to discover an approximation of a complex function with little effort. The main downside of this approach is the lack of guarantee regarding the convergence towards a sparse solution. The initial choice of the model’s coefficients appears to have a significant impact on the final solution.

### v2

@@ 3. The Nested SINDy approach
where $\mathcal{N}$ denotes our nested model, parameterized by $\theta$. Consequently, the resolution of this nonlinear optimization problem may be significantly more challenging and algorithms that work well in the linear case do not necessarily have the same performance in the nonlinear case.
The method then primarily depends on the choice of the architecture $\mathcal{N}$. The goal is to introduce new layers that achieve favorable trade-offs between expressivity and optimization complexity. In this work, we propose two architectures, which are described in the following sections: the PR model (in Section 3.2) and the PRP model (in Section 3.3). Both of are based on the polynomial layer, described in Section 3.1. From now on, we call the basic SINDy layer, given by a projection (2.2) onto $L(\mathcal{F})$, the radial layer.
3.1. Polynomial layers
To improve the expressivity of the usual SINDy approach, we introduce an additional layer, to be applied after the usual SINDy expression. This layer constructs a variety of monomials from the input variable $x$, represented as follows:
@@ 3.1. Polynomial layers
with $\omega_{0,0}$ acting as the constant term of the polynomial, effectively substituting the bias. Another choice is to limit the sum over $i+j$ to a maximum value $d$ to reduce the number of terms in the polynomial, and thus obtain bivariate polynomials up to degree $d$. This strategy is applied in Section 3.3; with two variables $x$ and $y$, such a polynomial layer reads:
$f_{\text{poly}}(x,y)=\sum_{0\leq i+j\leq d}\omega_{i,j}x^{i}y^{j}.$ (3.1)
3.2. The PR Model
The PR (Polynomial-Radial) model augments the basic SINDy framework by introducing a polynomial layer that operates before the usual SINDy radial layer. As an example, for one input variable, if =$\mathcal{F}=\{\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits,\cos\}$ and $d=2$, then the PR model can learn all functions with expression
@@ 3.2. The PR Model
This PR layer can also be seen as a pure polynomial layer combined with a linear layer. We, hence, consider the PR model to have four layers: a polynomial layer from Section 3.1, a (fully-connected) linear layer, a radial layer, and a final linear layer. The last two layers are identical to the standard SINDy model, while the first two layers are new additions. Figure 1 illustrates this structure.
@@
$f_{\theta,\text{PR}}(x)=\sum\limits_{j=1}^{l}c_{j}f_{j}\left(\sum\limits_{i=0}^{d}\omega_{i,j}x^{i}\right)+\beta,$ (3.2)
where $\theta$ includes all the trainable parameters of the model. The functions $f_{j}$ are derived from the specified function set $\mathcal{F}$, while $\omega_{i,j}$ are the weights of the polynomial layer, and $c_{j}$ and $\beta$ are the weights and bias of the final linear layer, respectively. The parameters $\omega_{i,j}$ now depend on $j$ since they are in the $j$-th radial layer. It is important to note that the radial and polynomial layers are not associated with any adjustable parameters.
@@ 3.3. The PRP Model
The PRP (Polynomial-Radial-Polynomial) model enhances the PR model by introducing an additional polynomial layer following the radial layer. This structure significantly improves the ability of the model to represent complex interactions within datasets. For example, the PRP model can express the function $f(x)=\arctan(x)\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x)$, assuming arctangent and sine are in the dictionary, while the PR model cannot, as it is not a linear combination of the functions contained within the dictionary. In the PRP model, the outputs of the radial layer are first processed through a fully-connected linear layer. This linear layer serves as an intermediate stage, transforming the outputs of the radial layer into a new set of variables. These variables are then fed into a subsequent polynomial layer, which allows for the formation of various monomial combinations of the outputs of the linear layers, such as squaring or multiplying them.
@@
$f_{\theta,\text{PRP}}(x)=\sum\limits_{1\leq\lvert i\rvert\leq d}\omega_{i_{1},i_{2},\dots,i_{l}}^{\text{PR}}f_{1,\text{PR}}(x)^{i_{1}}f_{2,\text{PR}}(x)^{i_{2}}\dots f_{k,\text{PR}}(x)^{i_{k}}+\beta^{\prime},$ (3.3)
where $\lvert i\rvert=i_{1}+i_{2}+\ldots+i_{k}$ is the length of the multi-index $(i_{1},\ldots,i_{k})$, $\theta$ includes all the trainable parameters of the model, $\omega_{i_{1},i_{2},\dots,i_{k}}^{\text{PR}}$ are the weights of the final linear layer, $\beta^{\prime}$ is the bias of the final linear layer, $k$ is the size of the output chosen for the intermediate linear layer (set to $k=2$ in our experiments), and $f_{1,\text{PR}}(x)$, $f_{2,\text{PR}}(x)$, …, $f_{k,\text{PR}}(x)$ are given by the following equation:
$f_{{k^{\prime}},\text{PR}}(x)=\sum\limits_{j=1}^{l}c_{j,k^{\prime}}f_{j}\left(\sum\limits_{i=0}^{d}\omega_{i,j}x^{i}\right)+\beta_{k^{\prime}}\text{, \quad forall }1\leq k^{\prime}\leq k.$
Figure 2 shows a graphical representation of this model.
@@
Figure 2. Structure of the PRP model for $k=2$, $d=2$, and 2 input dimensions.
@@ 4.1. Adding a regularization term to enforce sparsity
$\lambda(\text{epoch})=\lambda_{0}(1+0.5\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(\text{epoch}/10)),$ (4.1)
where $\lambda_{0}$ is the initial Lasso coefficient, and where the sine function is used to make the coefficient oscillate between $0.5\lambda_{0}$ and $1.5\lambda_{0}$.
@@ 4.4. Adding noise to the gradient of the loss function
The noise amplitude has to be well-tuned. It depends on the learning rate $\ell$ of the optimizer, the values of the parameters $\theta$, and on the actual mean squared error $L_{\text{MSE}}$:
$\nabla_{\theta}L\leftarrow\nabla_{\theta}L+\varepsilon(\ell,L_{\text{MSE}},\theta).$ (4.2)
with
$\varepsilon\sim\mathcal{N}\left(0,\alpha\times\ell\times\min(L_{\text{MSE}},1)\times\lvert\theta\lvert\right),\text{ and $\alpha$ a parameter to tune.}$.
@@ 5.1.1. Case 1: trigonometric function involving composition using the PR block
The hyperparameters and training settings are as follows:
  • Optimizer: the Adam optimization algorithm is used, with a learning rate of $10^{-3}$.
  • Weight initialization: the P and R layers do not use any weight initialization. The linear layer, however, applies a normal distribution initialization, where the weights are drawn from a Gaussian distribution with a specified mean ($\mu$) set to 0 and a standard deviation ($\sigma$) set to 1.5.
  • Batch: the training is performed on 10 0001000010\,00010 000 uniformly distributed values, with a batch size of $n_{\text{batch}}=1\,000$, resulting in 10 iterations per epoch.
  • Sparsity: Lasso regularization is applied at a rate of $10^{-3}$, and the model weights undergo pruning every $30$ epochs with a threshold set to $0.05$.
The function uses a patience counter to keep track of how many consecutive epochs have passed without significant improvement in the loss function. This is calculated by comparing the relative difference in loss between the current and previous epochs to a predefined threshold (here, $10^{-2}$). If the relative difference is lower than this threshold for $50$ consecutive epochs ($\texttt{patience}=50$), the training stops. In this case, the training stopped after $300$ epochs out of a maximum of 1 00010001\,0001 000.
The learning process of the PR model with the specified parameters successfully approximated the target function with the expression $x\mapsto-\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}-1.57)$ that captures the underlying behaviour of the target function within the considered domain, see Figure 3. Indeed, at the end of the training, the absolute and relative MSEs are respectively given by
$\int_{0}^{3}\left(-\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}-1.57)+\cos(x^{2})\right)^{2}\mathrm{d}x=8.31\times 10^{-7}\text{\quad and \quad}\frac{\int_{0}^{3}\left(\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}-1.57)+\cos(x^{2})\right)^{2}\mathrm{d}x}{\int_{0}^{3}cos^{2}(x^{2})\mathrm{d}x}=4.92\times 10^{-7}.$
We clearly see that the behaviour of the function is captured up to (single-precision) machine error. The sparsity of the solution is significant, with only 3 nonzero coefficients out of the original 29 parameters, underscoring the effectiveness of the model in identifying the essential structure of the function within the given domain.
@@
(d) Epoch 180: the model has approximated the target function.
Figure 3. Case 1 from Section 5.1.1: evolution of the learned function $x\mapsto\cos(x^{2})$ over successive training epochs. Each subfigure represents the PR model’s predictions (in blue) against the target function (in orange) at epochs 10, 70, 120, and 180, showcasing the model’s progressive learning and convergence towards approximation of the target function. We observe that the first epochs learn on the interval $[0,3]$, while subsequent epochs are able to generalize.
@@ 5.1.2. Case 2: trigonometric function involving composition using the PRP block
To further illustrate the capabilities of the introduced PRP-Nested SINDy method (3.3) described in Section 3.3, we examine the same function as in Section 5.1.1, namely $f:x\mapsto\cos(x^{2})$, considered over the interval $[0,3]$. This time, employing a dataset comprising 1 00010001\,0001 000 data points, the PRP-Nested SINDy model is tasked with replicating this function.
The model’s architecture begins with a polynomial (P) layer that computes monomials up to the second degree (i.e. with monomials $x$ and $x^{2}$). A subsequent linear layer then transforms these two polynomial features into nine outputs, incorporating a bias term in the process. The radial (R) layer follows, applying a suite of predefined functions to each node, belonging to the following dictionary:
@@
The architecture then introduces an intermediate linear layer with output size set to $k=2$, a second polynomial (P) layer still with $d=2$, i.e. with monomials $x$, $y$, $xy$, $x^{2}$, $y^{2}$, and a final linear layer, thus completing the structure of the model.
The hyperparameters match those of Section 5.1.1, except that the Adam learning rate is set to $10^{-4}$. We note that the loss function stagnates after around $300$ epochs.
The PRP model’s learning process successfully approximated the expression $x\mapsto\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}+1.57)$. This expression effectively captures the underlying behaviour within the specified domain (Figure 4). After training, the absolute and relative MSEs are respectively given by
$\int_{0}^{3}\left(\cos(x^{2})-\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}+1.57)\right)^{2}\mathrm{d}x=8.31\times 10^{-7}\text{\quad and \quad}\frac{\int_{0}^{3}\left(\cos(x^{2})-\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}+1.57)\right)^{2}\mathrm{d}x}{\int_{0}^{3}\cos^{2}(x^{2})\mathrm{d}x}=4.92\times 10^{-7}.$
Just like before, the behaviour of the function is captured up to (single-precision) machine error. The resulting model exhibits notable sparsity, with only 3 nonzero coefficients from the initial 31 parameters.
Figure 4. Case 2 from Section 5.1.2: comparison of the results from the PRP-Nested SINDy, which yielded the function $x\mapsto\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x^{2}+1.57)$, with the target function $x\mapsto\cos(x^{2})$, over the interval [0, 3].
@@ 5.1.3. Case 3: Trigonometric Function Multiplication using the PRP Block
The training parameters for learning the function $(x,y)\mapsto 2\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x)\cos(y)$ are specifically configured as follows:
  • Optimizer: the Adam optimization algorithm is employed with a base learning rate of $10^{-4}$, and is run for 1 00010001\,0001 000 epochs.
  • Weight initialization: in the uniform weight initialization for the linear layers, the standard deviation is set to $0.5$, and the mean is maintained at 00.
  • Batch: the training is performed on 10 0001000010\,00010 000 uniformly distributed values, with a batch size of $n_{\text{batch}}=1\,000$, resulting in 10 iterations per epoch.
  • Sparsity: the Lasso regularization rate is set to $0.1$, and the pruning of model weights occurs every $50$ epochs with an initial pruning at $10$ epochs, using a pruning threshold of $0.01$.
@@
with sparsity accounted for by 14 nonzero coefficients out of the original 32 parameters. A comparison of this function with the target function is illustrated in Figure 5. The MSE between the target function and its approximation is $5.69\times 10^{-2}$.
@@ 5.1.4. Case 4: Perimeter of an ellipse
Since any rescaled ellipse remains an ellipse, we assume without loss of generality that $b=1$ for the rest of this experiment. To assess the relevance of our Nested SINDy technique, we will compare its performance against the two reference solutions: Ramanujan’s approximation, and a linear interpolation between the two endpoints of the interval $[1,30]$. Indeed, using a linear interpolation makes sense for large values of $a$, since $P$ is equivalent to $4a$ when $a\to\infty$, as can be seen by inspecting (5.2).
To find formulas that approximate the perimeter of an ellipse, we train two PRP models. The first one is trained on the interval $[1,5]$ and the second one on the interval $[1,25]$. We run 50 learning sessions for each model and keep the model with the shortest associated formula. The only difference between the learning sessions is the seed used for initializing the random number generator. A learning session consists of two trainings, each with its own parameters:
  • Optimizer: the Adam optimization algorithm is used for both trainings. The first training is run for $100$ epochs with a base learning rate of $10^{-1}$, while the second one includes 1 50015001\,5001 500 epochs and a learning rate of $10^{-3}$.
  • Weight initialization: the initial weights are uniformly sampled with mean 00 and standard deviation $0.5$.
  • Batch: the training is performed on 1 00010001\,0001 000 uniformly distributed values, using a single batch.
  • Sparsity: the Lasso regularization rate is set to $10^{-1}$ in the first training and $10^{-2}$ in the second one. The pruning occurs after $30$ epochs below the threshold of $5\times 10^{-2}$ in both trainings.
Overall, the first training aims to move faster near a solution, while the second one aims to converge to a precise solution.
In its best session, the first model found a basic quadratic polynomial:
@@
The second model found a more complex expression which is too long to be displayed here. In this second model, 25 out of the 64 parameters are nonzero, corresponding to approximately 40% sparsity. The MSE obtained at the end of the training is $1.97$. This metric, as the previous ones, is biased because the training is stopped at an arbitrary step and there is no final tuning step. Hence, we perform a final tuning step, by writing the final expression of the model and performing a gradient descent to only minimize the MSE. The later mentioned results include this final tuning step and show that the effective MSE can be significantly reduced.
@@
Overall, this experiment provides an insight into the ability of the Nested SINDy approach to discover an approximation on a classical problem without knowledge other than the data points. It demonstrates that this method can be used to discover an approximation of a complex function with little effort. The main downside of this approach is the lack of guarantee to find a sparse solution. The initial choice of the model’s coefficients appears to have a significant impact on the final solution.